OpenStudy (beccaboo333):

When using the substitution method, what expression for x in the first equation would be substituted in the second equation? x + 3y = -7 2x + 4y = -4

OpenStudy (beccaboo333):

please help. medals and fanning will be rewarded

OpenStudy (beccaboo333):

@Kfins99

OpenStudy (anonymous):

x + 3y = -7 2x + 4y = -4 x= -7 -3y 2(-7 -3y)=-4

OpenStudy (anonymous):

\[y= \frac{ -10 }{ 2}\]

OpenStudy (anonymous):

@carlyn101